Solving Rational Equations.
1+(12)/(x^2-4)=(3)/(x-2)

Answer by stanbon(57307)

1+(12)/(x^2-4)=(3)/(x-2)
----
Multiply thru by (x^2-4) to get:
(x^2-4) + 12 = 3(x+2)
----
x^2-3x -4+12-6 = 0
--------
x^2 - 3x +2 = 0
---
Factor:
(x-2)(x-1) = 0
x = 2 or x = 1
==================
Cheers,
Stan H.
==================